# Phase separation of chemokinetic active particles

### v1

@@ Abstract
Motility-induced phase separation (MIPS) is a well-studied nonequilibrium collective phenomenon observed in active particles. Recently, there has been growing interest in how coupling the self-propulsion of active particles to chemical degrees of freedom affects MIPS. Previous studies have shown that incorporating chemotaxis and the production or consumption of chemicals by active particles results in various pattern formations, such as arrested phase separation and traveling waves. In this study, we demonstrate that similar phenomena can be induced when active particles consume chemicals and exhibit chemokinesis, where higher chemical concentrations enhance self-propulsion without causing alignment with the chemical gradient. We discover that MIPS is intensified if chemical consumption is proportional to particle density but is suppressed if chemical consumption is closely tied to particle motion. This leads to a wider range of collective behaviors, including arrested phase separation and oscillating patterns. Our findings are based on a hydrodynamic theory derived from a particle-based model via standard methods.
I Introduction
@@
Active matter, or substances consisting of self-propelled units, has attracted much attention as a novel class of nonequilibrium systems exhibiting various collective phenomena Ramaswamy (2010); Marchetti et al. (2013); Ramaswamy ; Jülicher et al. (2018); Gompper et al. (2020); Bowick et al. (2022); te Vrugt and Wittkowski . Among these is the motility-induced phase separation (MIPS), which refers to the spontaneous separation between high- and low-density regions stemming from the self-propulsion and short-range repulsion of the particles Tailleur and Cates (2008); Fily and Marchetti (2012); Redner et al. (2013); Bialké et al. (2013); Cates and Tailleur (2015). To describe the phenomena, various theoretical frameworks have been developed, including top-down derivations of active field theories Wittkowski et al. (2014); Tiribocchi et al. (2015); Tjhung et al. (2018) and bottom-up derivations of coarse-grained hydrodynamic models Bertin et al. (2006); Bialké et al. (2013). Predictions of these theories have also been tested by extensive numerics Stenhammar et al. (2013, 2014); Caporusso et al. (2020); Shi et al. (2020).
@@
For simplicity, these studies have focused on active particles that propel themselves at constant speed and change directions only on their own accord. However, there have also been studies of how the coupling between active particles and chemicals affects the collective phenomena. They have particularly focused on chemotaxis, i.e., the tendency of particles to move up (chemoattraction) or down (chemorepulsion) the chemical concentration gradient. If chemicals are produced by the particles themselves (autophoresis), it is known that both chemoattraction and chemorepulsion can induce phase separation Keller and Segel (1970, 1971). In particular, autophoresis with chemorepulsion leads to a variety of phase-separated patterns, such as dynamic clustering and traveling waves Liebchen et al. (2015, 2017). Furthermore, when chemicals are injected into the system at a constant rate and chemoattractants are taken up by particles, then the MIPS can be arrested with dynamic instabilities Zhao et al. (2023). These results indicate that the interplay between active particles and chemicals can induce nontrivial changes in the collective behaviors.
On the other hand, even in the absence of chemotaxis, the chemical concentration may only affect the motility of active particles. This phenomenon, called chemokinesis, is found in various artificial and biological systems. For example, it is known that the velocity of a Pt-coated Janus particle increases with the local hydrogen peroxide concentration Howse et al. (2007). Another example is given by cilia and flagella of motile bacteria, whose beat frequency is strongly affected by the surrounding ATP concentration Chen et al. (2015). When such chemokinetic effects are combined with the inhomogeneous concentration fields caused by the chemical consumption of active particles, one can naturally expect the emergence of nontrivial clustering patterns.
@@
To address this simple and practically relevant scenario, we explore a model system in which active particles are subject to both chemokinesis and chemical consumption. Through a hydrodynamic description of the system and linear stability analysis, we can analytically predict under which conditions the MIPS is enhanced or suppressed. We also observe microphase separation and oscillating patterns in our model, showing that chemokinetic effects are sufficient to induce a rich range of pattern formation behaviors.
The rest of the paper is organized as follows. In Sec. II, we introduce a model of chemical-consuming ABPs that exhibit chemokinesis, which is then coarse-grained to yield a hydrodynamic model. In Sec. III, we investigate phases and patterns generated by the hydrodynamic model, comparing predictions of the linear stability analysis with simulation results. In Sec. IV, the behaviors of the hydrodynamic model are compared with those of particle-based models. Finally, we summarize and conclude in Sec. V.
Figure 1: (a) Schematic illustrations of two different regimes of local chemical consumption. Phase diagram of the hydrodynamic model in (b) the BMR with $\lambda=0.01$, $D_{\mathrm{c}}=1$ and (c) the AMR with $\lambda=0.5$, $D_{\mathrm{c}}=0.1$. The LSA predictions for the phase boundaries are indicated by the dotted (H–S boundary), the dash-dotted (H–O boundary), and the dashed (S–O boundary) lines. For comparison, snapshots of particle distribution (with darker regions indicating higher density) obtained by hydrodynamic simulations are shown together. The values of $\lambda$ and $D_{\mathrm{c}}$ for each snapshot correspond to the coordinates of its center. The unphysical region with $v_{\mathrm{eff}}<0$ is grayed out. For this and all the other results in this paper, we use $\alpha=1$, $\rho_{0}=0.5$, $D=1$, $D_{\mathrm{r}}=3$, and $\kappa=20$, unless mentioned otherwise.
II Model
@@
$\dot{\mathbf{r}}_{k}(t)$ $=\mu\tilde{v}(n(\mathbf{r}_{k},t))\,\mathbf{e}_{k}-\sum_{k^{\prime}\neq k}\mu\bm{\nabla}_{k}V(|\mathbf{r}_{k}-\mathbf{r}_{k^{\prime}}|)$
@@
where $\mathbf{r}_{k}$ denotes the position of the $k$th ABP, $\mu$ the mobility of each ABP, $\tilde{v}$ the magnitude of self-propulsion, $\mathbf{e}_{k}$ the direction of self-propulsion exhibiting angular Brownian motion characterized by the rotational diffusion coefficient $D_{\mathrm{r}}$, $V$ the pairwise interaction between ABPs, $T$ the temperature of the medium, and $\bm{\xi}_{k}$ a Gaussian white noise satisfying $\langle{\bm{\xi}}_{k}(t)\rangle=0$ and $\langle{\bm{\xi}}_{k}(t){\bm{\xi}}_{k^{\prime}}(t^{\prime})\rangle=\delta_{kk^{\prime}}\mathbb{I}\delta(t-t^{\prime})$ for the $2\times 2$ identity matrix $\mathbb{I}$. Note that chemokinesis is incorporated into the model via the dependence of $\tilde{v}$ on the local chemical concentration $n(\mathbf{r}_{k},t)$. Unlike chemotaxis, which aligns or anti-aligns $\mathbf{e}_{k}$ to the concentration gradient, chemokinesis modifies only the motility of the active particle, as captured by the function $\tilde{v}(n)$. For simplicity, we focus on the linear regime where $\tilde{v}(n)=\alpha n$, with $\alpha$ controlling the strength of chemokinesis  [^1] The chemokinetic effect directly captured by this model is called orthokinesis. We note that there is also klinokinesis, which couples $D_{\mathrm{r}}$ to the chemical concentration. While the distinction between these two is important in the biological context, they are not fundamentally different in the physical context since a klinokinetic system can be mapped to an orthokinetic system by suitable rescaling of time. Thus, here we focus only on the latter..
@@
where $D_{\mathrm{c}}$ is the diffusion constant of the chemical, $\lambda$ is the consumption rate, and $f$ is a function governing how the chemical consumption depends on the state of all ABPs and the chemical. The behaviors of $f$ have been empirically studied for various species under different physical conditions Videler and Nolet (1990); Ohlberger et al. (2007). For simplicity, here we focus on two regimes of $f$ that would differently affect the phase separation behavior of ABPs, as schematically illustrated in Fig. 1(a).
@@
In the BMR, the chemical consumption rate increases with the density of ABPs, which would lower the chemical concentration within ABP clusters. The resulting chemokinetic slowdown hinders the ABPs from escaping the cluster. Thus, one can naturally expect that the phase-separating tendencies are strengthened in the BMR.
(ii) Active Metabolic Regime (AMR). In this regime, the chemical is mostly used when the ABPs actually move. Then the chemical consumption rate depends on how quickly the ABPs move in the self-propulsion direction, as expressed by
@@
In the AMR, the chemical consumption rate drops if the ABPs are stuck together, resulting in the accumulation of chemical within the ABP clusters. Then the ensuing chemokinetic speedup makes the cluster more prone to disintegration. Thus, in contrast to the BMR, one can expect that the phase-separating tendencies are weakened in the AMR.
The particle-based model described thus far can be coarse-grained following the standard procedure given in Ref. Bialké et al. (2013). This yields a system of hydrodynamic equations in terms of the particle density $\rho(\mathbf{r},t)$ and the chemical concentration $n(\mathbf{r},t)$, as shown below:
@@
so that the positive coefficient $\zeta$ quantifies the slowdown effect. We propose this hydrodynamic model, expressed by Eqs. (II) and (6), as a minimal model for exploring the effects of chemokinesis and chemical consumption on the MIPS, which is discussed in the next section.
III Linear stability analysis and hydrodynamic simulation
In this section, we investigate various phases of the hydrodynamic model expressed by Eqs. (II) and (6). Towards this end, we first conduct linear stability analysis (LSA) to predict when and how homogeneous solutions become unstable. Then, the analytical results are compared with numerical simulations of the hydrodynamic model.
Before moving on to the details, let us give a brief sketch of how LSA is done. The homogeneous density and concentration profiles $\rho(\mathbf{r},t)=\rho_{0}$ and $n(\mathbf{r},t)=n_{0}$ can solve Eq. (II) when the injection rate is given by
@@ III Linear stability analysis and hydrodynamic simulation
Assuming this condition, we let the density and the concentration profiles fluctuate by $\delta\rho(\mathbf{r},t)$ and $\delta n(\mathbf{r},t)$, respectively. Taking the Ansatz $\delta\rho=\delta\rho_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$ and $\delta n=\delta n_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$, Eq. (II) can be linearized as an eigenvalue equation
@@
See Appendix A for explicit expressions of the matrix elements $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$. The eigenvalue $\omega$ thus satisfies the characteristic equation $\omega^{2}-(\mathcal{A}+\mathcal{D})\omega+\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=0$, whose roots are given by
@@
Depending on how the real and the imaginary parts of $\omega_{\pm}$ vary with $q^{2}\equiv\mathbf{q}\cdot\mathbf{q}$, we can identify three phases. First, when $\mathrm{Re}\,\omega_{\pm}<0$ for all $q^{2}$, the homogeneous profiles are stable against fluctuations, resulting in the homogeneous (H) phase. Second, if $\mathrm{Re}\,\omega_{\pm}>0$ for some interval of $q^{2}$ and $\mathrm{Im}\,\omega_{\pm}=0$ there, then any slight deviations from the homogeneous profiles would grow exponentially until the LSA breaks down. Assuming that the lack of oscillation persists beyond the LSA regime, this behavior would result in a coexistence phase where high-density and low-density regions form a largely stable interface, which we call the stationary (S) phase. Finally, when there is an interval of $q^{2}$ with $\mathrm{Re}\,\omega_{\pm}>0$ and $\mathrm{Im}\,\omega_{\pm}\neq 0$, perturbations of the homogeneous profiles exhibit oscillations with growing amplitudes. We call this regime the oscillatory (O) phase.
While we have defined these three phases in terms of the behaviors of the system in the LSA regime, the true steady-state behaviors, which may lie well beyond the LSA regime, can be different. In the following, we separately discuss the LSA predictions of the three phases for the BMR and the AMR, which are then compared with numerics.
Figure 2: Evolution of particle density $\rho(\mathbf{r},t)$ and chemical concentration $n(\mathbf{r},t)$ in the hydrodynamic simulation of the BMR for (a) $n_{0}=18$, $\zeta=6$ and (b) $n_{0}=6$, $\zeta=6$. The other parameters are identical to those used in Fig. 1(b).
III.1 BMR facilitates phase separation
For the BMR, one can easily check that $\mathcal{B}\mathcal{C}>0$ for all $q^{2}$ as long as $v_{0}\equiv\alpha n_{0}-\rho_{0}\zeta>0$, which must be true in the regime where Eq. (6) is a good approximation. Thus, the BMR lacks the O phase, and we only need to consider the boundary between the H and the S phases. As detailed in Appendix A, the boundary is derived as
@@ III.1 BMR facilitates phase separation
When $\lambda=0$, there is no chemical consumption, so $n(\mathbf{r},t)$ stays at its initial value $n_{0}$. Since the chemical concentration does not change anywhere, the effective self-propulsion strength can be written as $v(\rho,n)=v_{0}-\rho\zeta$, where $v_{0}\equiv\alpha n_{0}$ is a fixed constant. In this case, chemokinesis becomes completely irrelevant, effectively reducing the dynamics of the system to those of the vanilla model decoupled from the chemicals. Then, as detailed in Appendix A, the phase boundary is determined by the effective diffusion coefficient
@@
that governs the collective relaxation of density profiles when only the short-range repulsion between particles is taken into account (via $\zeta$). The MIPS occurs when and only when $D_{\text{eff}}<0$, which yields the phase boundary
$D_{\text{vanilla}}=\frac{(\alpha n_{0}-\rho_{0}\zeta)(2\rho_{0}\zeta-\alpha n_{0})}{2D_{\mathrm{r}}}\;$ (12)
@@
so that $D>D_{\text{vanilla}}$ ($D<D_{\text{vanilla}}$) corresponds to the H phase (S phase). Comparing the above formula with Eq. (10), it is clear that $D_{\text{BMR}}>D_{\text{vanilla}}$ always holds, indicating that nonzero $\lambda$ broadens the S phase at the expense of the H phase. Thus, as previously expected, the BMR facilitates MIPS.

### v2

@@ Abstract
Motility-induced phase separation (MIPS) is a well-studied nonequilibrium collective phenomenon observed in active particles. Recently, there has been growing interest in how coupling the self-propulsion of active particles to chemical degrees of freedom affects MIPS. Although the effects of chemotaxis on MIPS have been extensively studied, little is known about how chemokinesis affects MIPS. In this study, we demonstrate that various patterns can be induced when active particles consume chemicals and exhibit chemokinesis, where higher chemical concentrations enhance self-propulsion without causing alignment with the chemical gradient. We discover that MIPS is intensified if chemical consumption is proportional to particle density (as in the basal metabolic regime), but it is suppressed if chemical consumption is closely tied to particle motion (as in the active metabolic regime). While the former produces large-scale phase separation via coarsening, the latter suppresses the coarsening process, leading to microphase separation and oscillating patterns. We also derive a hydrodynamic theory that describes these findings.
Introduction
@@
Active matter, or substances consisting of self-propelled units, has attracted much attention as a novel class of nonequilibrium systems exhibiting various collective phenomena [1, 2, 3, 4, 5, 6, 7, 8]. Among these is the motility-induced phase separation (MIPS), which refers to the spontaneous separation between high- and low-density regions stemming from self-propulsion and short-range repulsion of active particles [9, 10, 11, 12, 13]. Various theoretical frameworks have been developed to describe the phenomena, including top-down derivations of active field theories [14, 15, 16] and bottom-up derivations of coarse-grained hydrodynamic models [17, 12]. Predictions of these theories have been tested by extensive simulations [18, 19, 20, 21].
@@
For simplicity, these studies have focused on active particles that propel themselves at constant force and change directions only on their own accord. However, there have also been studies of how the coupling between active particles and chemicals affects the collective phenomena. They have mainly focused on chemotaxis, i.e., the tendency of microorganisms or particles to move up (chemoattraction) or down (chemorepulsion) the chemical concentration gradient. At the hydrodynamic level, such effects have been modeled as the parallel or antiparallel alignment of local flux [22, 23, 24, 25] or polarization [26, 27] with the concentration gradient. If the chemicals are released by the particles themselves (autochemotaxis), it is well known that chemoattraction leads to complete phase separation [22, 23, 24], while chemorepulsion generates a variety of patterns, such as finite-size dynamic clusters and traveling waves [26, 27]. Moreover, when externally injected chemoattractants are taken up by active particles, MIPS can be arrested or create oscillating patterns [25]. On the other hand, when chemotaxis arises as a consequence of the diffusiophoresis of spherical particles, the coupling between active particles and chemicals has been shown to induce both complete phase separation and dynamic clustering, where multiple small clusters coexist with particles constantly joining and leaving them [28, 29, 30]. Dynamic clustering of the platinum–gold Janus particles, empirically observed in [31], was attributed to this effect. These results demonstrate that the interplay between active particles and chemicals can induce a host of pattern-formation behaviors.
However, there is another type of coupling between active particles and chemicals. For example, while the direction of self-propulsion is unaffected by the chemical concentration gradient, the magnitude of self-propulsion may change depending on the concentration level. This phenomenon, called chemokinesis, occurs in various kinds of cells including sperm cells [32, 33], neural cells [34], leukocytes [35, 36], and bacteria [37, 38, 39, 40, 41] as they respond to external chemical stimuli. While chemokinesis often coexists with chemotaxis in microorganisms, different types of receptors are responsible for chemokinetic and chemotactic responses, implying that the strengths of those responses can be controlled separately [42]. Indeed, there are both naturally occurring and artificially induced examples of cells which exhibit strong chemokinesis but very little chemotaxis, such as neutrophils in aged human subjects [43] and human endothelial cells after the CMG2 gene is blocked [44]. There are also synthetic active particles known to exhibit chemokinetic response to their fuel with no or little chemotaxis. These include platinum–polystyrene Janus particles propelled by diffusiophoresis [45] and platinum–gold rods propelled by electrophoresis [46], which are both fueled by hydrogen peroxide in water. For the Janus particles driven by diffusiophoresis, analyses show that chemotaxis is negligible when the mobility constants of the two halves of the particle are not very different [28, 29, 47].
Then, how does chemokinesis affect the pattern formation of active matter? The question has been addressed mainly in the context of how active particles respond to an externally imposed chemical concentration gradient. For example, a model study predicted that chemokinesis mixed with chemotaxis enhances the migratory response of bacteria to an imposed chemoattractant concentration gradient compared to those only capable of chemotaxis [48]. In experiments with self-propelled rods with an externally imposed fuel concentration gradient, chemokinesis was found to be responsible for the accumulation of particles in the low-concentration region [46]. Broadening the scope to include light-activated particles, analogous effects can also be found in photokinetic bacteria, which have been used to control their self-organized patterns and transport [49, 50, 51, 52].
@@
But what if the active particles create the concentration gradients as they take up the chemicals? Little has been studied about how chemokinesis in the absence of chemotaxis affects the phase-separation and pattern-formation behaviors of autonomous systems formed by active particles and chemicals. To address this scenario, we explore a model system in which active particles are subject to chemokinesis and chemical consumption. We observe that chemokinetic effects are sufficient to induce a rich range of pattern formation behaviors, which can be described analytically through a hydrodynamic model and linear stability analysis.
Figure 1: Schematic illustrations of two different regimes of local chemical consumption. Each blue elliptical particle with a tail represents an active particle, whose direction of self-propulsion is indicated by a green arrow. Regions colored in light blue correspond to active particle clusters. Chemicals fueling the active particles are represented by filled orange dots, while by-products produced by chemical consumption are depicted with empty orange dots. (a) In the Basal Metabolic Regime, where chemical consumption rate is proportional to particle density, chemical concentration tends to be higher where particle density is lower. This enhances phase separation, creating a single large-scale cluster. (b) In the Active Metabolic Regime, where chemical consumption rate is determined by particle flux, chemical concentration tends to be lower where particle density is higher. This suppresses phase separation, creating multiple finite-size clusters.
Methods
Particle-based model
@@ Methods
$\dot{\mathbf{r}}_{k}(t)$ $=\tilde{v}(n(\mathbf{r}_{k},t))\,\mathbf{e}_{k}-\sum_{k^{\prime}\neq k}\mu\bm{\nabla}_{k}V(|\mathbf{r}_{k}-\mathbf{r}_{k^{\prime}}|)$
@@
where $\mathbf{r}_{k}$ denotes the position of the $k$th ABP, $\tilde{v}$ the magnitude of self-propulsion, $\mu$ the mobility of each ABP, $\mathbf{e}_{k}$ the direction of self-propulsion exhibiting angular Brownian motion characterized by the rotational diffusion coefficient $D_{\mathrm{r}}$, $V$ the pairwise interaction between ABPs, $T$ the temperature of the medium, and $\bm{\xi}_{k}$ a Gaussian white noise satisfying $\langle{\bm{\xi}}_{k}(t)\rangle=0$ and $\langle{\bm{\xi}}_{k}(t){\bm{\xi}}_{k^{\prime}}^{\textsf{T}}(t^{\prime})\rangle=\delta_{kk^{\prime}}\mathbb{I}\delta(t-t^{\prime})$ for the $2\times 2$ identity matrix $\mathbb{I}$. Note that chemokinesis is incorporated into the model via the dependence of $\tilde{v}$ on the local chemical concentration $n(\mathbf{r}_{k},t)$. Unlike chemotaxis, which aligns or anti-aligns $\mathbf{e}_{k}$ to the concentration gradient, chemokinesis modifies only the motility of the active particle, as captured by the function $\tilde{v}(n)$. For simplicity, we focus on the linear regime where $\tilde{v}(n)=\alpha n$, with $\alpha>0$ controlling the strength of chemokinesis.
The chemokinetic effect directly captured by this model is called orthokinesis. We note that there is also klinokinesis, which couples $D_{\mathrm{r}}$ to the chemical concentration. While the distinction between these two is crucial in the biological context, they are not fundamentally different in the physical context since a klinokinetic system can be mapped to an orthokinetic system by suitable rescaling of time. Thus, here we focus only on the latter.
@@
where $D_{\mathrm{c}}$ is the diffusion constant of the chemical, $\lambda$ is the consumption rate, and $f$ is a function governing how the chemical consumption depends on the state of all ABPs and the chemical. If $\lambda=0$, the particle dynamics is decoupled from the chemical, which we call the vanilla model. The behaviors of $f$ have been empirically studied for various species under different physical conditions [53, 54]. For simplicity, here we focus on two regimes of $f$ that would differently affect the phase separation behavior of ABPs.
@@
In the BMR, the chemical consumption rate increases with the density of ABPs, which would lower the chemical concentration within ABP clusters. The resulting chemokinetic slowdown hinders the ABPs from escaping the cluster. Thus, one can naturally expect that the phase-separating tendencies are strengthened in the BMR, as schematically illustrated in Fig. 1(a).
(ii) Active Metabolic Regime (AMR). In this regime, the chemical consumption is tightly coupled to the motion of the ABPs, as expressed by
@@
Note that the chemical is consumed when the self-propulsion orientation $\mathbf{e}_{k}$ and the velocity $\dot{\mathbf{r}}_{k}$ are in the same direction, while any motion of the particle against the self-propulsion leads to the chemical production. This is a natural consequence of the mechanochemical coupling obeying the Onsager reciprocal relations, as discussed in [4, 55, 56, 57, 58, 59].
In the AMR, the chemical consumption rate depends on both the particle density and the effective speed along the self-propulsion direction. If the particle density is sufficiently high within the ABP clusters formed by MIPS, the decrease of particle speed overwhelms the increase of particle density, reducing the chemical consumption and raising the chemical concentration within the ABP clusters. Then the ensuing chemokinetic speedup makes the cluster more prone to disintegration. Thus, in contrast to the BMR, one can expect that the phase-separating tendencies may be weakened in the AMR, as schematically illustrated in Fig. 1(b).
To investigate the effects of chemical consumption and chemokinesis on phase separation and pattern formation, we directly simulate the particle-based model following Eqs. (Particle-based model)–(4), with the short-range repulsion between particles given by the Weeks–Chandler–Andersen (WCA) potential
where $u_{0}$ sets the magnitude of the potential, and $\sigma$ is the effective diameter of each particle. We fix $u_{0}=100$ and $\sigma=1$, which, together with the persistence time fixed at $\tau_{\mathrm{p}}\equiv D_{\mathrm{r}}^{-1}=1/3$, define the natural units of length, time, and energy, respectively. We note that the Dirac delta function appearing in Eqs. (3) and (4) are discretized in the simulation, so that its value becomes $1$ only when the particle is in the grid corresponding to $\mathbf{r}$ and 00 otherwise. The simulation results are discussed in the Results section.
Figure 2: Phase separation behaviors of the particle-based model with repulsive interactions. (a) Fraction of particles belonging to the largest cluster (identified using the DBSCAN algorithm [60]) as the global chemical concentration $n_{0}$ is varied. Error bars indicate standard deviations obtained using six different realizations. Insets: Snapshots displaying the locations of particles (blue dots) and the heatmaps of chemical concentration (red shades), which are sampled from the four selected data points of the main figure. Note that chemical concentration is irrelevant and thus not shown for the vanilla model. The system is a two-dimensional torus of side length $L=200$ and global particle density $\rho_{0}=1/2$. We also use $\alpha=1$, $D_{0}\equiv\mu T=1$, $D_{\mathrm{r}}=3$, $\sigma=1$, $u_{0}=100$, $\lambda=0.1$, and $D_{\mathrm{c}}=1$. See the discussions below Eqs. (Particle-based model), (2), and (5) for descriptions of the parameters. We also show the histograms of local particle density as $n_{0}$ is varied for (b) the Basal Metabolic Regime (BMR), (c) the vanilla model, and (d) the Active Metabolic Regie (AMR). All physical quantities and snapshots are measured at time $t=200$. See Supplementary Movie 1A (1B) for an animated comparison between the vanilla model and the BMR (AMR) at $n_{0}=20$ ($n_{0}=30$).
Hydrodynamic model
To understand the phase separation and pattern formation behaviors exhibited by the particle-based model, we also propose a system of hydrodynamic equations for the particle density $\rho(\mathbf{r},t)$ and the chemical concentration $n(\mathbf{r},t)$, which can be derived from the particle-based model assuming slow spatial variations of the fields and fast angular relaxation of the ABPs. See Appendix A for the detailed derivation, which closely follows the procedure given in Ref. [12]. The resulting hydrodynamic equations are given by
@@
so that the positive coefficient $\zeta$ quantifies the slowdown effect. We propose the hydrodynamic model expressed by Eqs. (Hydrodynamic model) and (7) as a minimal model for exploring the effects of chemokinesis and chemical consumption on MIPS.
Results
Behaviors of the particle-based model
To verify whether the BMR (AMR) facilitates (suppresses) MIPS, we let the particle-based model described in the Methods section evolve from a homogeneous initial state to $t=200$ and observe its phase separation behaviors. In Fig. 2(a), we show the fraction of particles belonging to the largest cluster (also called the giant cluster) for the BMR, the AMR, and the vanilla model using the same set of parameters (except for $\lambda=0$ used in the vanilla model). The BMR develops a larger particle cluster compared to the vanilla model, with the chemical concentration becoming higher outside the particle cluster. In contrast, the AMR strongly suppresses the growth of clusters, with the chemical concentration becoming higher inside the particle cluster. We highlight that the giant clusters formed by the BMR have smooth interfaces, while those formed by the AMR tends to be of rugged shapes. This rules out the presence of the Ostwald process driven by the positive interfacial tension in the AMR, which is crucial for the large-scale phase separation in the BMR.
The same conclusion can also be drawn from the histograms of the local particle density for the BMR [see Fig. 2(b)], the vanilla model [see Fig. 2(c)], and the AMR [see Fig. 2(d)] at different values of $n_{0}$. For the BMR, the separation between dense and sparse regions occurs at the smallest value of $n_{0}$ among the three cases, and the density gap at a fixed value of $n_{0}$ is also the largest. For the AMR, on the contrary, the phase separation can hardly be detected by the histograms. In other words, the AMR does not exhibit phase separation, or the dense particle clusters are too small to be captured by the histograms.
These results indicate that the BMR facilitates the large-scale MIPS, whereas the AMR suppresses MIPS and allows only microphase separation. As discussed earlier, the latter effect is attributable to increased self-propulsion of the ABPs inside the high-density particle clusters. In more realistic situations, the self-propulsion of ABPs may eventually saturate to an upper bound instead of growing linearly with the chemical concentration as in Eq. (7). In this case, we observe that the suppression of MIPS in the AMR becomes weaker, with the particle clusters becoming larger as the upper bound on self-propulsion becomes tighter. This lends further support to our view that increased self-propulsion suppresses MIPS. See Appendix C and Fig. A2 for more detailed discussions.
Behaviors of the hydrodynamic model
Linear stability analysis
The hydrodynamic model expressed by Eqs. (Hydrodynamic model) and (7) allows us to predict the phase-transition behaviors of the system through linear stability analysis (LSA) of the homogeneous solution. We begin by noting that the homogeneous density and concentration profiles $\rho(\mathbf{r},t)=\rho_{0}$ and $n(\mathbf{r},t)=n_{0}$ can solve Eq. (Hydrodynamic model) when the injection rate is given by
@@ Results
Assuming this condition, we let the density and the concentration profiles fluctuate by $\delta\rho(\mathbf{r},t)$ and $\delta n(\mathbf{r},t)$, respectively. Taking the Ansatz $\delta\rho=\delta\rho_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$ and $\delta n=\delta n_{0}e^{i\mathbf{q}\cdot\mathbf{r}+\omega t}$, Eq. (Hydrodynamic model) can be linearized as an eigenvalue equation
@@
See Appendix B for explicit expressions of the matrix elements $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$. The eigenvalue $\omega$ thus satisfies the characteristic equation $\omega^{2}-(\mathcal{A}+\mathcal{D})\omega+\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}=0$, whose roots are given by
@@
Depending on how the real and the imaginary parts of $\omega_{\pm}$ vary with $q^{2}\equiv\mathbf{q}\cdot\mathbf{q}$, we can identify three phases. First, when $\mathrm{Re}\,\omega_{\pm}<0$ for all $q^{2}$, the homogeneous profiles are stable against fluctuations, resulting in the homogeneous (H) phase. Second, if $\mathrm{Re}\,\omega_{\pm}>0$ for some interval of $q^{2}$ and $\mathrm{Im}\,\omega_{\pm}=0$ there, then any slight deviations from the homogeneous profiles would grow exponentially until the LSA breaks down. Assuming that the lack of oscillation persists beyond the LSA regime, this behavior would result in a coexistence phase where high-density and low-density regions form a largely stable interface, which we call the stationary (S) phase. Finally, when there is an interval of $q^{2}$ with $\mathrm{Re}\,\omega_{\pm}>0$ and $\mathrm{Im}\,\omega_{\pm}\neq 0$, perturbations of the homogeneous profiles exhibit oscillations with growing amplitudes. We call this regime the oscillatory (O) phase. See Fig. A1 and Appendix B for detailed derivations of the phase boundaries.
Figure 3: Snapshots and phase diagrams of the hydrodynamic model. These are shown for (a) the Basal Metabolic Regime with $\lambda=0.5$, $D_{\mathrm{c}}=0.1$ and for the Active Metabolic Regime with (b) $\lambda=0.01$, $D_{\mathrm{c}}=1$ and (c) $n_{0}=20$, $\zeta=22$. Note that $n_{0}$ denotes the global chemical concentration, $\lambda$ the chemical consumption rate, $D_{\mathrm{c}}$ the diffusion constant of the chemical, and $\zeta$ the coefficient describing the deceleration due to interparticle interactions. Predictions of linear stability analysis for the phase boundaries are indicated by the dotted (H–S boundary), the dash-dotted (H–O boundary), and the dashed (S–O boundary) lines. For comparison, snapshots of particle distribution (with darker regions indicating higher density) obtained by hydrodynamic simulations are shown together. The values of $\lambda$ and $D_{\mathrm{c}}$ for (a) and (b), and $\lambda$ and $D_{\mathrm{c}}$ for (c) in each snapshot correspond to the coordinates of its center. The unphysical region with $v_{\mathrm{eff}}\equiv\alpha n_{0}-\rho_{0}\zeta<0$ is grayed out for (a) and (b). For this and all the following results in this paper, we use $\alpha=1$, $\rho_{0}=1/2$, $D=1$, $D_{\mathrm{r}}=3$, and $\kappa=20$, unless mentioned otherwise. See Supplementary Movies 2A, 2B, and 2C for animated versions of (a), (b), and (c), respectively.
Now, based on the LSA described so far, we investigate various phases of the hydrodynamic model. In the following, we separately discuss the LSA predictions for the phase boundaries in the BMR and in the AMR, which are then compared with simulation results.
BMR facilitates phase separation
For the BMR, one can easily check that $\mathcal{B}\mathcal{C}>0$ for all $q^{2}$ as long as $v_{\mathrm{eff}}\equiv\alpha n_{0}-\rho_{0}\zeta>0$, which must be true in the regime where Eq. (7) is a good approximation. Thus, the BMR lacks the O phase, and we only need to consider the boundary between the H and the S phases. As detailed in Appendix B, the boundary is derived as
@@
When $\lambda=0$, there is no chemical consumption, so $n(\mathbf{r},t)$ stays at its initial value $n_{0}$. Since the chemical concentration does not change anywhere, the effective self-propulsion strength can be written as $v(\rho,n)=v_{0}-\rho\zeta$, where $v_{0}\equiv\alpha n_{0}$ is a fixed constant. In this case, chemokinesis becomes completely irrelevant, effectively reducing the dynamics of the system to those of this vanilla model decoupled from the chemicals. Then, as detailed in Appendix B, the phase boundary is determined by the effective diffusion coefficient
@@
that governs the collective relaxation of density profiles when the self-propulsion and the short-range repulsion between particles are taken into account (via $v_{0}$ and $\zeta$, respectively). MIPS occurs when and only when sufficiently strong deceleration effects (represented by $\zeta$) induce $D_{\text{eff}}<0$, so that ‘negative diffusion’ amplifies the density fluctuations. This yields the phase boundary
$D_{\text{vanilla}}=\frac{(v_{0}-\rho_{0}\zeta)(2\rho_{0}\zeta-v_{0})}{2D_{\mathrm{r}}}\;,$ (13)
@@
with $D>D_{\text{vanilla}}$ ($D<D_{\text{vanilla}}$) corresponding to the H phase (S phase). Comparing the above formula with Eq. (11), it is clear that $D_{\text{BMR}}>D_{\text{vanilla}}$ always holds, indicating that nonzero $\lambda$ broadens the S phase at the expense of the H phase. Thus, as previously expected, the BMR facilitates MIPS.